# PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models

### v3

@@ 2.3 Belief State Space
Different from the perfect observation case, the update of the belief state is nonlinear for the general observation model. This yields much difficulty for us to use value functions and dynamic programming methods to analyze the problem. Given an initial belief state, it appears that the size of the belief state space will grow exponentially over time as all possible realizations of observations/rewards/feedback are incessantly traversed and updated. Fortunately, a large number of numerical experiments have shown that in the sense of Euclidean norm approximation, the state space exhibits convergence after a process of iterative calculations (see Sec. 3 for details). Note that the initial belief state is $\boldsymbol{\omega}$, $\{\mathcal{B}_{l}\}_{0\leq l\leq L}$ is a list of operators defined by feedback states and state update rules, that is, for $1\leq l\leq L$, $\mathcal{B}_{l}\boldsymbol{\omega}$ is the updated state caused by the $l$-th feedback state under the active action, and $\mathcal{B}_{0}$ is the state update operator under the passive action. In this paper, $\mathcal{B}_{0}\boldsymbol{\omega}=\boldsymbol{\omega}P$. Define the $T$-step state space recursively as follows:
@@
$\Omega(T|\boldsymbol{\omega})=\begin{cases}\{\mathcal{B}_{l_{1}}\cdots\mathcal{B}_{l_{T}}\boldsymbol{\omega}:0\leq l_{1},\cdots,l_{T}\leq L\}\cup\Omega(T-1|\boldsymbol{\omega}),&T\geq 2\\ \{\mathcal{B}_{l}\boldsymbol{\omega}:0\leq l\leq L\}\cup\{\boldsymbol{\omega}\},&T=1\end{cases}.$
@@
Under the definition of the $T$-step state space, $\Omega(+\infty|\boldsymbol{\omega_{0}})$ is the belief space obtained by traversing all possible state update rules starting from the initial state $\boldsymbol{\omega_{0}}$. Clearly, $\Omega(+\infty|\boldsymbol{\omega_{0}})$ is countable and we denote it by $\Omega(\boldsymbol{\omega_{0}})$. Under a policy $\pi$, let $p_{ij}(\pi):=P(\boldsymbol{\omega}(t+1)=\boldsymbol{\omega_{j}}|\boldsymbol{\omega}(t)=\boldsymbol{\omega_{i}})$ be the conditional probability of transition from belief state $\boldsymbol{\omega_{i}}$ to $\boldsymbol{\omega_{j}}$. Due to different actions of activation and passivity affecting the state update, decomposition of $p_{ij}(\pi)$ should be carried out. Let $a(t)$ be the indicator function of whether the arm is activated in slot $t$, then we define the transition probability of the belief state as follows:
@@
By (14)-(15) and the total probability theorem, we have
$p_{ij}(\pi)$ $=\sum\limits_{a=0}^{1}P(a(t)=a|\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i})p_{ij}^{a}.$ (16)
The elements in the probability transition matrix under passive and active actions have the following more specific expressions:
@@
=$=\sum\limits_{\{l:\mathcal{B}_{l}\boldsymbol{\omega}_{i}=\boldsymbol{\omega}_{j},1\leq l\leq L\}}P(F(t)=l|\text{active and }\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i}).$
@@ 2.4 Two-Arm Problem and Achievable Region
With the rapid development at the junction of operations research, stochastic optimization and reinforcement learning, quite a few effective methods for finite-state problems have been found, among which achievable region with conservation laws is a great example. Bertsimas and Niño-Mora (1996) and Niño-Mora (2001) adopted the analytical framework of generalized conservation laws and partial conservation laws for the classic MAB and RMAB problems with finite state spaces, respectively, and provided efficient index algorithms for the corresponding problems. In the rest subsections, we will apply the PCL framework to the analysis of our RMAB problem with an infinite state space, and theoretically build the foundation for the construction of an efficient index policy.
@@
$I_{\boldsymbol{\omega}}^{1}(t)=\begin{cases}1,&\text{if the arm is active in slot $t$ and its belief state is $\boldsymbol{\omega}$}\\ 0,&\text{otherwise}\end{cases}$
@@
$I_{\boldsymbol{\omega}}^{0}(t)=\begin{cases}1,&\text{if the arm is passive in slot $t$ and its belief state is $\boldsymbol{\omega}$}\\ 0,&\text{otherwise}\end{cases}.$
@@
$x_{\omega}^{a}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{+\infty}\beta^{t}I_{\boldsymbol{\omega}}^{a}(t)\right],\quad a=0,1$ (19)
$x_{\omega}^{a}(\pi,\boldsymbol{\omega}_{i})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{+\infty}\beta^{t}I_{\boldsymbol{\omega}}^{a}(t)\bigg{|}\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad a=0,1$ (20)
where $\boldsymbol{\omega}(t)$ is the belief state of the arm in slot $t$. In Whittle’s relaxation, the Lagrangian multiplier $\lambda$ can be regarded as a subsidy when the arm is made passive. Thus the optimization objective of the partially observable RMAB problem with subsidy can be written as follows:
$\max_{\pi}$ $\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}+\lambda\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega}}^{0}$ (21)
subject to $x_{\boldsymbol{\omega}_{i}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0}=e_{\boldsymbol{\omega}_{i}}+\beta(\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega}_{j}}^{1}+\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega}_{j}}^{0}),\quad\forall\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$
@@
where $R_{\boldsymbol{\omega}}=\sum\limits_{i=1}^{M}\omega_{i}R_{i},\boldsymbol{\omega}=(\omega_{1},\cdots,\omega_{M})$, $p_{ij}^{a}$ are given by (17)-(18), $e_{\boldsymbol{\omega}_{i}}$ is the indicator whether the initial belief state is $\boldsymbol{\omega}_{i}$, $r_{i}$ is the expected immediate reward when $S(t)=i$ with the active action. Obviously, $R_{\boldsymbol{\omega}}$ is bounded above and we can assume $|R_{\boldsymbol{\omega}}|\leq C,\forall\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})$. For the equality constraints of this optimization problem, it can be understood that the left side of the equation is the direct performance measure for state $\boldsymbol{\omega}_{i}$, while the right side of the equation is another way to represent it. That is, each occurrence of state $\boldsymbol{\omega}_{i}$ in the system is transitioned by another state $\boldsymbol{\omega}_{j}$. Therefore, the performance measure for state $\boldsymbol{\omega}_{i}$ can be represented by a combination of performance measures of other states. This model with subsidy can be explained more intuitively through a two-arm system. In this system, the first arm is the original arm while the second arm (auxiliary arm) has only one state 0. In every slot we choose one of the two arms to activate. When the auxiliary arm is activated, we obtain a fixed reward $\lambda$. Our goal is to find a policy to maximize the long-term expected discounted reward by deciding which arm to choose in each slot. The objective function can thus be written as
$\max\limits_{\pi}\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}(\pi)+\lambda x_{0}^{1}(\pi),$
@@ 2.5 Partial Conservation Law
In this section, we elaborate the relationship between performance measures for each state in the two-arm system. To measure the performance of the original arm under policy $\pi$, we define the following variables:
@@
For the auxiliary arm, we have $A_{0}^{\{0\}}=1$ and $W_{0}^{\{0\}}=\lambda$. Generally, we can extend these variables to the multi-arm case. Assume that there are two arms and their state spaces are $\Omega(\boldsymbol{\omega})$ and $\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ with $\Omega(\boldsymbol{\omega})\cap\Omega^{\prime}(\boldsymbol{\omega}^{\prime})=\emptyset$. Then for $\Omega\subset\Omega(\boldsymbol{\omega})$, $\Omega^{\prime}\subset\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ and $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega})$, we define $T_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=T_{\boldsymbol{\omega_{i}}}^{\Omega},R_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=R_{\boldsymbol{\omega_{i}}}^{\Omega},A_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=A_{\boldsymbol{\omega_{i}}}^{\Omega}$ and $W_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=W_{\boldsymbol{\omega_{i}}}^{\Omega}$. Previously, Niño-Mora provided the relationships among $T_{\boldsymbol{\omega_{i}}}^{\Omega}$, $R_{\boldsymbol{\omega_{i}}}^{\Omega}$, $A_{\boldsymbol{\omega_{i}}}^{\Omega}$, and $W_{\boldsymbol{\omega_{i}}}^{\Omega}$ for finite-state problems in (Niño-Mora, 2001, 2002). Here, we prove that these conservation relations also hold for countable state spaces:
@@
Assume there exists a constant $C$, such that $|R_{\boldsymbol{\omega_{i}}}|\leq C$ holds for any $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ and $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, the following two identities hold:
@@
Since $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0})=\frac{1}{1-\beta}$, and for any $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$, $x_{\boldsymbol{\omega_{i}}}^{1},x_{\boldsymbol{\omega_{i}}}^{0}\geq 0$, thus by the Cauchy criterion for series convergence, $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}$ and $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{0}$ are both absolutely convergent. Due to the countability of $\Omega(\boldsymbol{\omega_{0}})$, summing the above constraint equations gives:
@@
Therefore, $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}$ and $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0}$ are both convergent and their summation orders are exchangeable. By constraint equations and convergence of series, we have
@@
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}$
@@
The equality holds when policy $\pi$ gives higher priority to state 0 (the second arm) than states in $\Omega^{c}$. However, not all $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$ have such property, so we need to consider a family of sets $\tilde{\Omega}\subset 2^{\Omega(\boldsymbol{\omega_{0}})}$, which satisfies $\Omega(\boldsymbol{\omega_{0}})\in\tilde{\Omega}$. For any $\Omega\in\tilde{\Omega}$, define $b(\Omega):=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\},b(\{0\}\cup\Omega)=\inf\limits_{\pi}\{x_{0}^{1}(\pi)+\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$. Here we give the following definition of partial conservation law generalized from Niño-Mora (2001):
@@
The problem is said to satisfy the partial conservation law with respect to $\tilde{\Omega}$ if, for any $\Omega\subset\tilde{\Omega}$ and policy $\pi$, the following hold:
@@
We call collection family $C=\{\Omega_{\boldsymbol{\omega_{i}}}^{C}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\subset\tilde{\Omega}$ is a full subset chain of $\tilde{\Omega}$, if it satisfies the following conditions:

### v4

@@ 2.3 Belief State Space
Different from the perfect observation case, the update of the belief state is nonlinear for the general observation model. This yields much difficulty for us to use value functions and dynamic programming methods to analyze the problem. Given an initial belief state, it appears that the size of the belief state space will grow exponentially over time as all possible realizations of observations/rewards/feedback are traversed and updated. Fortunately, a large number of numerical experiments have shown that in the sense of Euclidean norm approximation, the approximate state space exhibits stability after a process of iterative calculations (see Sec. 3 for details). Note that the initial belief state is $\boldsymbol{\omega}$, $\{\mathcal{B}_{h}\}$ is a complete list of operators defined by observed and feedback states and state update rules. Set $H:=|\{\mathcal{B}_{h}\}|$. Define the $T$-step state space recursively as follows:
@@
$\Omega(T|\boldsymbol{\omega})=\begin{cases}\{\mathcal{B}_{h_{1}}\cdots\mathcal{B}_{h_{T}}\boldsymbol{\omega}:1\leq h_{1},\cdots,h_{T}\leq H\}\cup\Omega(T-1|\boldsymbol{\omega}),&T\geq 2\\ \{\mathcal{B}_{h}\boldsymbol{\omega}:1\leq h\leq H\}\cup\{\boldsymbol{\omega}\},&T=1\end{cases}.$
@@
Under the definition of the $T$-step state space, $\Omega(+\infty|\boldsymbol{\omega_{0}})$ is the belief space obtained by traversing all possible state update rules starting from the initial state $\boldsymbol{\omega_{0}}$. Clearly, $\Omega(+\infty|\boldsymbol{\omega_{0}})$ is countable (since the belief update happens only at discrete time indices, see Sec. 2.1) and we denote it by $\Omega(\boldsymbol{\omega_{0}})$. Under a policy $\pi$, let $p_{ij}(\pi):=P(\boldsymbol{\omega}(t+1)=\boldsymbol{\omega_{j}}|\boldsymbol{\omega}(t)=\boldsymbol{\omega_{i}})$ be the conditional probability of transition from belief state $\boldsymbol{\omega_{i}}$ to $\boldsymbol{\omega_{j}}$. Due to different actions of activation and passivity affecting the state update, decomposition of $p_{ij}(\pi)$ should be carried out. Let $a(t)$ be the indicator function of whether the arm is activated in epoch $t$, then we define the transition probability of the belief state as follows:
@@
Clearly, $p_{ij}^{0}$ and $p_{ij}^{1}$ do not depend on $\pi$ since the action is already specified in the condition. By (14)-(15) and the law of total probability, we have
$p_{ij}(\pi)$ $=\sum\limits_{a=0}^{1}P_{\pi}(a(t)=a|\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i})p_{ij}^{a}.$ (16)
Note that $P_{\pi}(a(t)=a|\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i})$ depends on $\pi$ since the policy specifies which action to take given the current belief state. The elements in the probability transition matrix under passive and active actions have the following more specific expressions:
@@
$=\sum\limits_{\{h:\mathcal{B}_{h}\boldsymbol{\omega}_{i}=\boldsymbol{\omega}_{j},1\leq h\leq H\}}\mathcal{B}_{h}\boldsymbol{\omega}_{i}P(h\text{-th belief update rule}|\text{active and }\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i}).$
@@ 2.4 Two-Arm Problem and Achievable Region
With the rapid development at the junction of operations research, stochastic optimization and reinforcement learning, quite a few effective methods for finite-state problems have been found, among which achievable region with conservation laws is a great example. bertsimas1996conservation and nino2001restless adopted the analytical framework of generalized conservation laws and partial conservation laws for the classic MAB and RMAB problems with finite state spaces, respectively, and provided efficient index algorithms for the corresponding problems. In the rest of the subsections, we will apply the PCL framework to the analysis of our RMAB problem with an infinite state space, and theoretically build the foundation for the construction of an efficient index policy.
@@
[…]
@@
[…]
@@
$x_{\boldsymbol{\omega}}^{a}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{+\infty}\beta^{t}I_{\boldsymbol{\omega}}^{a}(t)\right],\quad a=0,1$ (19)
$x_{\boldsymbol{\omega}}^{a}(\pi,\boldsymbol{\omega_{i}})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{+\infty}\beta^{t}I_{\boldsymbol{\omega}}^{a}(t)\bigg|\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad a=0,1$ (20)
where $\boldsymbol{\omega}(t)$ is the belief state of the arm in epoch $t$. In Whittle’s relaxation, the Lagrangian multiplier $\lambda$ can be regarded as a subsidy when the arm is made passive. Thus the optimization objective of the partially observable RMAB problem with subsidy can be written as follows:
$\max$ $\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}+\lambda\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega}}^{0}$ (21)
subject to $x_{\boldsymbol{\omega}_{i}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0}=e_{\boldsymbol{\omega}_{i}}+\beta\left(\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega}_{j}}^{1}+\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega}_{j}}^{0}\right),\quad\forall\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$
@@
where $R_{\boldsymbol{\omega}}=\sum\limits_{i=1}^{M}\omega_{i}R_{i},\boldsymbol{\omega}=(\omega_{1},\cdots,\omega_{M})$, $p_{ij}^{a}$ are given by (17)-(18), $e_{\boldsymbol{\omega}_{i}}$ is the indicator whether the initial belief state is $\boldsymbol{\omega}_{i}$, $R_{i}$ is the known expected immediate reward conditional on $S(t)=i$ with the active action. Note that $R_{\boldsymbol{\omega}}$ is uniformly bounded over $\Omega(\boldsymbol{\omega_{0}})$ since $|R_{\boldsymbol{\omega}}|\leq\max\{|R_{i}|\}$. As we will see in Sec. 2.5, both $\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega}}^{1}$ and $\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega}}^{0}$ are absolutely convergent so the objective function in (21) is well defined. The $\max$ operator is justified by observing that the objective function is upper bounded by $\frac{\max\{|R_{i}|\}+|\lambda|}{1-\beta}$ and our following analysis aims at finding the optimal stationary policy $\pi$ to maximize the objective function (i.e., the expected total discounted reward). Together with the constraint on each variable, our formulation falls into the standard form in  anderson1987linear. We point out the LP problem with a countably infinite number of both variables and constraints is significantly more complex than a finite-dimensional one and we must use the infinite-dimensional LP theory for investigating the duality and optimality, as detailed below.
For the equality constraints of this optimization problem, it can be understood that the left side of the equation is the direct performance measure for state $\boldsymbol{\omega}_{i}$, while the right side of the equation is a decomposition of its occupied time over previous states transitioning into it. Therefore, the performance measure for state $\boldsymbol{\omega}_{i}$ can be represented by a combination of performance measures of other states. This model with subsidy can be explained more intuitively through a two-arm system. In this system, the first arm is the original arm while the second arm (auxiliary arm) has only one state 0. In every epoch we choose one of the two arms to activate. When the auxiliary arm is activated, we obtain a fixed reward $\lambda$. Our goal is to find a policy to maximize the long-term expected discounted reward by deciding which arm to choose in each epoch. The benefit of using the two-arm model for the single-arm bandit with subsidy is that the second arm never changes its state and only the selected arm yields reward. Specifically, the objective function can be simply written as
$\max\limits\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}(\pi)+\lambda x_{0}^{1}(\pi),$
@@ 2.5 Partial Conservation Law
In this section, we elaborate the relationship between performance measures for each state in the two-arm system. We first introduce some important variables following a similar line of nino2001restless; nino2002dynamic. To measure the performance of the original arm under policy $\pi$, we define the following variables:
@@
For the auxiliary arm, we have $A_{0}^{\{0\}}=1$ and $W_{0}^{\{0\}}=\lambda$. Generally, we can extend these variables to the multi-arm case. Assume that there are two arms and their state spaces are $\Omega(\boldsymbol{\omega})$ and $\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ with $\Omega(\boldsymbol{\omega})\cap\Omega^{\prime}(\boldsymbol{\omega}^{\prime})=\emptyset$. Then for $\Omega\subset\Omega(\boldsymbol{\omega})$, $\Omega^{\prime}\subset\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ and $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega})$, we define $T_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=T_{\boldsymbol{\omega_{i}}}^{\Omega},R_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=R_{\boldsymbol{\omega_{i}}}^{\Omega},A_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=A_{\boldsymbol{\omega_{i}}}^{\Omega}$ and $W_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=W_{\boldsymbol{\omega_{i}}}^{\Omega}$. Previously, Niño-Mora provided the relationships among $T_{\boldsymbol{\omega_{i}}}^{\Omega}$, $R_{\boldsymbol{\omega_{i}}}^{\Omega}$, $A_{\boldsymbol{\omega_{i}}}^{\Omega}$, and $W_{\boldsymbol{\omega_{i}}}^{\Omega}$ for finite-state problems in (nino2001restless; nino2002dynamic). After the above formulation, we prove the important result that these conservation relations also hold for general countable state spaces:
@@
For any $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ and $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, the following two identities hold:
@@
Since $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0})=\frac{1}{1-\beta}$, and for any $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$, $x_{\boldsymbol{\omega_{i}}}^{1},x_{\boldsymbol{\omega_{i}}}^{0}\geq 0$, thus by the bounded convergence theorem for monotonic sequences (Theorem 3.14 in rudin1976principles), $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}$ and $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{0}$ are both absolutely convergent. Due to the countability of $\Omega(\boldsymbol{\omega_{0}})$, summing the above constraint equations gives:
@@
Therefore, $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}$ and $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0}$ are both absolutely convergent by the bounded convergence theorem for monotonic sequences and their summation orders are exchangeable (Theorem 8.3 in rudin1976principles). By constraint equations and absolute convergence of series, we have
@@
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}$
@@
The equality holds when policy $\pi$ gives higher priority to state 0 (the second arm) than states in $\Omega^{c}$. However, not all $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$ have such property, so we need to consider a family of sets $\tilde{\Omega}\subset 2^{\Omega(\boldsymbol{\omega_{0}})}$, which satisfies $\Omega(\boldsymbol{\omega_{0}})\in\tilde{\Omega}$. For any $\Omega\in\tilde{\Omega}$, define $b(\Omega):=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\},b(\{0\}\cup\Omega)=\inf\limits_{\pi}\{x_{0}^{1}(\pi)+\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$. Here we give the following definition of partial conservation law generalized from nino2001restless:
@@
Let $\tilde{\Omega}$ be a feasible set family on the full belief-state space $\Omega(\boldsymbol{\omega}_{0})$ (i.e., $\emptyset,\ \Omega(\boldsymbol{\omega}_{0})\in\tilde{\Omega}$, and the complement of any element in $\tilde{\Omega}$ is taken with respect to $\Omega(\boldsymbol{\omega}_{0})$). The problem is said to satisfy the partial conservation law (PCL) with respect to $\tilde{\Omega}$ if, for any $\Omega\in\tilde{\Omega}$ and policy $\pi$, the following hold:
$A_{\boldsymbol{\omega}}^{\Omega}>0,~\forall\boldsymbol{\omega}\in\Omega,$
@@
We call a collection family $C=\{\Omega_{\boldsymbol{\omega_{i}}}^{C}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\subset\tilde{\Omega}$ a full subset chain of $\tilde{\Omega}$, if it satisfies the following conditions: